# SI_appendix_phys

### v1

@@ IV Traveling solutions in large systems
For $L\gg 1$, some active-matter and non-reciprocal models [9, 40] support traveling phase-separated states (“traveling bands”) where a large system has macroscopic liquid and vapor domains separated by sharp interfaces, which move at constant velocity. However, this is not possible in the APLG.
To see this, substitute (6) into (1) and write $z=x-ct/L$ to obtain
@@
where $\varrho_{\sigma}$ and $\mathrm{m}=\varrho_{+}-\varrho_{-}$ denote the densities and magnetization in the traveling frame [recall (6)], and primes indicate derivatives with respect to $z$. Within the bulk of the phases, $\varrho^{\prime}=0$ so $\mathrm{m}=0$ there. Summing (7) over $\sigma$ to obtain the total density $\varrho$ and integrating then yields $c\varrho+\varrho^{\prime}-\text{Pe}(1-\varrho)\mathrm{m}=J$ where $J$ is an integration constant. Evaluating this expression inside the two phases where $\mathrm{m}=0=\varrho^{\prime}$, one finds $\varrho=J/c$ so both phases would need to have equal densities, ruling out any traveling band states (the special case $c=0=J$ recovers the PS state). This exact analysis illustrates the value of our exact hydrodynamic description: it is very difficult to extrapolate such results for large systems based on numerical simulations alone, especially because time-stepping (1) is expensive in large domains.
Among T states that do exist, natural solutions of (7) have $\varrho^{\prime}_{\sigma}=O(1/L)$ [with $c=O(1)$ and $\mathrm{m}=O(1/L)$], so densities vary smoothly on the macroscopic scale. We do find such solutions, but Fig. 4(a) includes an interfacial region where $\varrho$ varies quickly in space, hinting at the existence of solutions with traveling narrow interfaces [$\varrho^{\prime}=O(1)$].
@@
These states do indeed exist. We find them by the (multi-scale) method of matched asymptotic expansions [43], with results shown in Fig. 5. Specifically, we seek a solution to (7) with a sharp interface at $z=0$, and we split the domain of $z$ into two overlapping subdomains, an inner region around the wave front where $|z|=O(1)$ and an outer region far from the front ($|z|\gg 1$). In each region, we solve simplified equations whose solutions are ‘stitched together’ again to obtain an accurate solution to (7). Details of the calculation are given in Methods. In the inner region, this recovers the same Eqs. (9,10,11) that govern PS solutions. That is, sharp interfaces in T states are the same as liquid-vapor interfaces in PS, and connect points on the binodal, connected by tie-lines (recall Fig. 1). In the outer region, we obtain
$-(c/L)\varrho_{\sigma}^{\prime}=\partial_{z}\left[d_{s}(\varrho)\varrho_{\sigma}^{\prime}+\varrho_{\sigma}{\cal D}(\varrho)\varrho^{\prime}\right]\\ +\frac{\text{Pe}^{2}}{2}\partial_{z}\left\{\left[\varrho_{\sigma}s(\varrho)+|\sigma|d_{s}(\varrho)\right]\partial_{z}\left[d_{s}(\varrho)\varrho_{a}\right]\right\},$ (8)
@@
which is to be solved with periodic boundary conditions at $z=\pm L/2$ and matching conditions to the binodal densities as $z\to 0^{\pm}$. We numerically solve (8) for the densities $\varrho_{\sigma}$ and the speed $c$. Fig. 5(b) and Fig. 4(a) contrast solutions for $L=O(1)$ and $L\gg 1$.
We identify different types of T solutions on varying the volume fractions $\phi_{p},\phi_{a}$. The top row of Fig. 5 shows four solutions obtained by solving (8), and the bottom row shows the same solutions overlaid on the phase diagram, together with the corresponding solutions to the full/original equation (7). There is a very good agreement between the asymptotic solutions and those in finite domains, confirming the validity of the matched asymptotic construction.
@@
For pairs $(\phi_{a},\phi_{p})$ in the upper part of region D, we observe smooth (periodic) T solutions with no inner region, see Fig. 5(a). Reducing $\phi_{p}$, in the phase diagram, we find solutions with a single interface (Fig. 5(b)). The inner region occurs between points P2 and P3 and lies along a tie line between two points on the binodal. Further reducing $\phi_{p}$, the high-density part of the solution enters the binodal, leading to T solutions with two interfaces. Fig. 5(c) shows this transition point: the first inner region is a single point (P1), tangent to the binodal at its critical value, and the second interface occurs between P2 and P3. Reducing $\phi_{p}$ still further, we obtain solutions with two interfaces (Fig. 5(d)): these have two inner regions that both follow tie-lines.
@@
The array of solutions in Fig. 5 illustrates the rich phenomenology of the APLG in large domains. In particular, the appearance of narrow interfaces in T solutions is a surprising feature, since it connects these dynamically-patterned states to the equilibrium-like constructions of the binodal and the associated interfacial profiles. The distinction between narrow interfaces and macroscopically smooth profiles would be very challenging to characterize by direct numerical solution of (1): the matched asymptotics method makes this possible.
V Multistability
@@
We have characterized T solutions, but this does not guarantee that a time-dependent system will converge to a T state, nor even that they are stable. To address this question, Fig. 6 shows the types of long-time solutions obtained by numerically time-stepping (1). The system is initialized with a zero-magnetization homogeneous state, perturbed by the most unstable eigenmode of the linear stability analysis. In the case of complex eigenvalues, solutions in Fig. 6(a) use a linear combination of left- and right-moving modes, while Fig. 6(b) uses only the left-moving mode. These systems converge to steady states, which we characterize as H, PS, T, or CP (see Methods for further details).
@@
The results are consistent with Fig. 1 and demonstrate the existence and stability of T and CP states, for suitable parameters. Comparing Figs. 6(a,b), the steady state also depends on the initial conditions: initializing with a left-moving mode favors T solutions while symmetric initialization favors CP. Fig. 6(b) also shows the range of parameters over which we were able to find T solutions via matched asymptotics, showing that non-symmetric initialization may not be sufficient to drive the system into T states, even if they exist. Together, these observations demonstrate multiple dynamical attractors, as may be expected for such complex PDE systems. Fig. S4 shows an explicit example where both T and PS solutions exist for the same parameters. Understanding the domains of attraction of different states in more detail is an important challenge for future work.
VI Discussion
@@
We introduced the APLG as a microscopic model of interacting particles and characterized its hydrodynamic behavior. In addition to PS states familiar from pure active systems, we find rich behavior, including that the spinodal curve can protrude through the binodal. This signals the existence of dynamical steady states, which we classify according to their symmetries. Some of these results are similar to previous work on the NRCH equation [10, 9, 11, 13], but our hydrodynamic PDE includes the magnetization $m$ as a slow hydrodynamic variable, in addition to the two conserved densities $\rho_{a},\rho_{0}$; it also has a distinct set of nonlinearities. Hence, our approach of deriving hydrodynamic equations exactly from a microscopic model complements the generic description of non-reciprocal systems via the NRCH equation.
@@
To tame the complexity of the APLG’s behavior, we focussed on large domains $L\gg 1$. This enables numerically exact computation of the binodal and spinodal curves and precise characterization of traveling solutions that involve sharp interfaces. Surprisingly, interfacial profiles in static and traveling states both obey (4), which also describes the tie-lines in the phase diagram. If such connections are generic in non-reciprocal systems, they would have broad consequences for understanding their phase diagrams, including possibilities for long-ranged order similar to equilibrium. We also demonstrated multistability: qualitatively different solutions can exist, for the same parameters.
@@
Our results also raise further questions for the APLG, including the existence of dynamical states in large systems with patterns on finite length scales, and associated questions of wavelength selection [13, 44, 11]. While this work analyzed deterministic hydrodynamic equations, the theory of fluctuating hydrodynamics can also be developed for such models [20, 45] by retaining corrections to the limit $h\to 0$. This enables studies of metastability and the role of noise in determining a system’s eventual steady state. There might also be interesting corrections to the large-$L$ behavior studied here. One may also expect new behavior on replacing the two-state active-particle orientation ($\sigma=\pm$) with a continuous degree of freedom: hydrodynamic limits can be derived in this case [21, 23] but the resulting equations are challenging to analyze. Future work should investigate these issues.
@@ Methods
For large systems, $L\gg 1$, stationary phase-separated (PS) states contain large domains of liquid and vapor phases. The phases’ (total) volume fractions are denoted $\phi_{l},\phi_{v}$, which we compute by generalising the method of [23, 46, 32]. Specifically, we use (1) to construct equations of motion for $\rho_{0}$ and $\rho$; setting $\partial_{t}\rho=0=\partial_{t}\rho_{0}$ and integrating yields
@@
In the bulk of either phase, we have $\partial_{x}\rho_{\sigma}=0$, so from (11) $m=0$ there. Then evaluating Eqs. (9,10) in the bulk shows that $J=0=J_{0}$. Using this fact and eliminating $m$ between Eqs. (9,10), one obtains using (2) that $\partial_{x}\log[\rho_{0}/(1-\rho)]=0$ so $\rho_{0}=\nu(1-\rho)$ for some constant $\nu$. Integrating this equation over the whole domain, one obtains the value $\nu$ in (5). Then, combining Eqs. (4,9,11), we obtain a condition on $\rho(x)$ alone:
@@
with $g_{0},\Lambda,\kappa$ given in Appendix B.
The method of [23, 46, 32] can now be applied directly to (12). Within the bulk of the coexisting phases one has $\partial_{x}\rho=0=\partial_{x}^{2}\rho$, showing that $g_{0}(\phi_{l})=g_{0}(\phi_{v})$. In addition, (12) can be used to construct an effective free energy $\Phi$, from which $\phi_{l},\phi_{v}$ can be fully determined by a common tangent construction, see Appendix B for details. The compositions of the phases are then given by (4). A numerical implementation of this procedure yields the binodal curves in Fig. 1.
@@
To analyse the stability of homogeneous stationary solutions of (1), we introduce a perturbation to $\rho_{\sigma}$ constant of the form $\delta A^{\sigma}\exp(\lambda t+iqxt)$ for $\delta\ll 1$, and $\lambda$ can be obtained as the eigenvalue of a $3\times 3$ matrix. If $\text{Re}(\lambda)>0$, then the perturbation grows, signaling that the homogeneous state is unstable. The spinodal is the boundary between the regions of stable and unstable homogeneous solutions. Full details are given in Appendix C.
Traveling solutions via the method of matched asymptotics
To analyze T solutions in large domains, we define $\epsilon=1/L$ and seek an asymptotic solution of (7) as $\epsilon\ll 1$ using the method of matched asymptotic expansions. Recall that $z\in[-L/2,L/2]$ and that, without loss of generality, the interface is centered at $z=0$ (see Fig. 5b). We assume there is only one interface; the generalization to multiple interfaces is straightforward. The outer region is $|z|\gg 1$: we set $\bar{z}=\epsilon z$ and define $\varrho(z)=\bar{\varrho}(\bar{z})$ and $\mathrm{m}(z)=\bar{\mathrm{m}}(\bar{z})$ so (7) becomes
@@
with periodic boundary conditions at $\bar{z}=\pm 1/2$.
Expanding $\bar{\varrho}_{\sigma}$ and $\bar{\mathrm{m}}$ in powers of $\epsilon$, $\bar{\varrho}_{\sigma}\sim\bar{\varrho}_{\sigma}^{(0)}(\bar{z})+\epsilon\bar{\varrho}_{\sigma}^{(1)}(\bar{z})+\cdots$ and $\bar{\mathrm{m}}\sim\bar{\mathrm{m}}^{(0)}+\epsilon\bar{\mathrm{m}}^{(1)}+\epsilon^{2}\bar{\mathrm{m}}^{(2)}+\cdots$, we find that the leading- and first-order of (14) lead to $\bar{\mathrm{m}}^{(0)}=0$ and $\bar{\mathrm{m}}^{(1)}=-\text{Pe}\partial_{\bar{z}}[d_{s}(\bar{\varrho}^{(0)})\bar{\varrho}_{1}^{(0)}]$, respectively. After eliminating $\bar{\mathrm{m}}^{(2)}$, the $O(\epsilon^{2})$ of (14) leads to (8) at leading order in $\epsilon$. To solve (8), it remains to determine the boundary conditions as we enter the inner region, $|\bar{z}|\to 0$.
For the inner region $z=O(1)$, define densities $\hat{\varrho}_{\sigma}(z)=\varrho_{\sigma}(z)$. They solve (7), together with the matching condition to the outer region, $\lim_{z\to\pm\infty}\hat{\varrho}(z)=\lim_{\bar{z}\to 0^{\pm}}\bar{\varrho}(\bar{z})$, these limits exist under the assumptions of a localized interface. As $L\to\infty$ the LHS of (7) vanishes: this ensures consistency with the argument above that traveling bands do not exist, and justifies the scaling of the speed as $c/L$ in (7). Hence, the leading-order inner solution solves same equations as the PS state [Eqs. (9,10,11) with $J=0=J_{0}$], and interfaces in T states connect points on the binodal curve.
@@
Particle model: The APLG is a continuous time Markov chain on a finite state space. We simulate it exactly with the Gillespie algorithm [36], initially placing a $\sigma$-particle on a lattice site with probability $\phi_{\sigma}$, where $\phi_{\pm}=\phi_{a}/2,\phi_{0}=\phi_{p}$. In Figs. 2,3, the simulated domain has $\ell_{y}=\ell_{x}/4$; we plot the $y-$averaged values of the mesoscopic densities (see (S4)) with radius $r=0.1$.
Time-stepping (1): We use a first-order finite-volume scheme in space and forward Euler with adaptive time-stepping in time to obtain time-dependent numerical solutions $\rho_{\sigma}(x,t)$ to the hydrodynamic PDE (1), building on the numerical scheme of [23, 39] (see Appendix 1).
Classification of the steady-states in Fig. 6. We classify solutions of (1) into Homogeneous (H), Phase-separated (PS), Traveling (T) or Counter-propagating (CP) uses the following two metrics: the approximate speed $\tilde{c}(t):=\|\partial_{t}\rho_{\sigma}\|_{2}/\|\partial_{x}\rho_{\sigma}\|_{2}$ and the $L_{2}$ distance from uniform
@@
We solve (1) until a final time $t^{*}\geq 700$ when one of the following conditions is satisfied:
@@
    $d_{\text{H}}(t^{*})\geq 0.05$ and $\sup_{t\in[t^{*}-500,t^{*}]}\tilde{c}(t)<0.01$to PS solution.
@@
    $d_{\text{H}}(t^{*})\geq 0.05$, $\tilde{c}(t^{*})\geq 0.01$ and $\sup_{t\in[t^{*}-500,t^{*}]}|\tilde{c}^{\prime}(t)|<10^{-5}$to T solution.

### v2

@@ Results
Recall that $L\gg 1$ corresponds to $\ell_{x}\gg\sqrt{D_{T}/D_{R}}$: this means that the system size is much larger than the intrinsic length scale associated with particle motion. Some active-matter and nonreciprocal models [9, 48] support traveling phase-separated states where a large system has macroscopic liquid and vapor domains separated by sharp interfaces, which move at constant velocity. However, this is not possible in the APLG. To see this, substitute (6) into (1) and write $z=x-ct/L$ to obtain
@@
where $\varrho_{\sigma}$ and $\mathrm{m}=\varrho_{+}-\varrho_{-}$ denote the densities and magnetization in the traveling frame [recall (6)], and primes indicate derivatives with respect to $z$. Within the bulk of the phases, $\varrho^{\prime}=0$ so $\mathrm{m}=0$ in the bulk. Summing (7) over $\sigma$ to obtain the total density $\varrho$ and integrating then yields $(c/L)\varrho+\varrho^{\prime}-\text{Pe}(1-\varrho)\mathrm{m}=J$ where $J$ is an integration constant. Evaluating this expression inside the two phases where $\mathrm{m}=0=\varrho^{\prime}$, one finds $\varrho=LJ/c$ so both phases would need to have equal densities, ruling out any traveling phase-separated states (the special case $c=0=J$ recovers the PS state). This exact analysis illustrates the value of our exact hydrodynamic description: it is very difficult to extrapolate such results for large systems based on numerical simulations alone, especially because time-stepping Eq. (1) is expensive in large domains.
There are macroscopically smooth T solutions of (7) satisfying $\varrho^{\prime}_{\sigma}=O(1/L)$ [with $c=O(1)$ and $\mathrm{m}=O(1/L)$]. We do find such solutions numerically (see Methods). However, Fig. 4(a) includes an interfacial region where $\varrho$ varies quickly in space, hinting at the existence of solutions with traveling narrow interfaces [$\varrho^{\prime}=O(1)$].
Figure 6: T solutions of the finite-$L$ problem (7) for different values of $L$ and the large-$L$ problem (8) at $\text{Pe}=7.5$. Total density $\varrho$ (top) and magnetisation $\mathrm{m}$ (bottom). The magnetisation for $L=\infty$ is $\mathrm{m}=0$ for $|z|>0$. Parameters as in Fig. 5(b).
@@
These T states with sharp interfaces do indeed exist. We find them systematically using the method of matched asymptotic expansions [49], with results shown in Fig. 5. This method consists of expanding solutions in an asymptotic series in the small parameter $1/L$: this is a controlled approximation scheme that provides accurate results for large systems. Specifically, we seek a solution to (7) with an interface at $z=0$, and we split the domain of $z$ into two overlapping subdomains, an inner region around the wave front where $|z|=O(1)$ and an outer region far from the front ($|z|\gg 1$). Details of the calculation are given in Methods. In the inner region, the problem reduces, at leading-order in $1/L$, to the same Eqs. (9,10,11) that govern PS solutions. That is, the leading order of sharp interfaces in T states coincide with liquid-vapor interfaces in PS, and connect points on the binodal by tie-lines (recall Fig. 1). In the outer region, we can eliminate $\mathrm{m}$, and the system reduces to two equations
$-(c/L)\varrho_{a}^{\prime}=$ $\partial_{z}\left[d_{s}(\varrho)\varrho_{a}^{\prime}+\varrho_{a}{\cal D}(\varrho)\varrho^{\prime}\right]$ (8a)
$+\frac{\text{Pe}^{2}}{2}\partial_{z}\left\{\left[\varrho_{a}s(\varrho)+d_{s}(\varrho)\right]\partial_{z}\left[d_{s}(\varrho)\varrho_{a}\right]\right\},$
$-(c/L)\varrho_{0}^{\prime}=$ $\partial_{z}\left[d_{s}(\varrho)\varrho_{0}^{\prime}+\varrho_{0}{\cal D}(\varrho)\varrho^{\prime}\right]$ (8b)
$+\frac{\text{Pe}^{2}}{2}\partial_{z}\left\{\varrho_{0}s(\varrho)\partial_{z}\left[d_{s}(\varrho)\varrho_{a}\right]\right\},$
@@
where $\varrho_{a}=\varrho_{+}+\varrho_{-}$. Eq. (8) is to be solved with periodic boundary conditions at $z=\pm L/2$ and matching conditions to the binodal densities as $z\to 0^{\pm}$.
We numerically solve (8) for the densities $\varrho_{a},\varrho_{0}$ and the speed $c$ and show the results for different combinations of $(\phi_{a},\phi_{p})$ in Fig. 5. The top row of Fig. 5 shows four solutions of (8), and the bottom row shows the same solutions overlaid on the phase diagram, together with the corresponding solutions to the original finite-$L$ problem (7). We see that the finite-$L$ solutions follow the tie lines, similar to the matched asymptotic solutions. Fig. 6 shows how direct solutions of (7) approach the asymptotic solution as $L$ is increased, confirming that the matched asymptotic analysis provides accurate results. Further, the spinodal curves of the outer problem (8) agree with the spinodals of the full problem (see Appendix 3).
It is in the outer region formulation (8) that the effective nonreciprocal interactions become most clearly apparent in the hydrodynamic equation (1). In particular, we show in Appendix B that (8) can be expressed as a nonlinear cross-diffusion system for the active and passive densities $\rho_{a}$, $\rho_{0}$ with nonreciprocal effective interactions between them.
@@
We identify different types of T solutions on varying the volume fractions $\phi_{p},\phi_{a}$. For pairs $(\phi_{a},\phi_{p})$ in the upper part of region D, we observe smooth (periodic) T solutions with no inner region, see Fig. 5(a). Reducing $\phi_{p}$ in the phase diagram, we find solutions with a single interface (Fig. 5(b)). The inner region occurs between points P1 and P2 and lies along a tie line between two points on the binodal. Another interesting case is when the high-density part of the solution enters the binodal, leading to T solutions with two interfaces. Fig. 5(c) shows this transition point: the first inner region is a single point (P1), tangent to the binodal at its critical value, and the second interface occurs between P2 and P3. Slightly reducing $\phi_{a}$, we obtain solutions with two interfaces (Fig. 5(d)): these have two inner regions that both follow tie-lines.
@@
The array of solutions in Fig. 5 illustrates the rich phenomenology of the APLG in large domains. In particular, the appearance of narrow interfaces in T solutions is a surprising feature, since it connects these dynamically-patterned states to the equilibrium-like constructions of the binodal and the associated interfacial profiles. While the analytical characterization of T solutions is not possible for CP solutions, numerical simulations show that CP solutions also feature narrow interfaces that follow the tie lines (see Appendix, Fig. S5). The distinction between narrow interfaces and macroscopically smooth profiles would be very challenging to characterize by direct numerical solution of (1): the method of matched asymptotics makes this possible.
.5 Multistability
@@
We have characterized T solutions, but this does not guarantee that the time-dependent system will converge to a T state, nor even that they are locally stable. There could also be other T solutions with different speeds $c$. To address this question, Fig. 7 shows the types of some long-time solutions obtained by numerically time-stepping (1). The system is initialized with a zero-magnetization homogeneous state, perturbed by the most unstable eigenmode of the linear stability analysis. In the case of complex eigenvalues, solutions in Fig. 7(a) use a linear combination of left- and right-moving modes, while Fig. 7(b) uses only the left-moving mode. These systems converge to steady states, which we characterize as H, PS, T, or CP (see Methods for further details).
@@
The results are consistent with Fig. 1 and demonstrate the existence and stability of T and CP states, for suitable parameters. Comparing Figs. 7(a,b), the steady state also depends on the initial conditions: initializing with a left-moving mode favors T solutions while symmetric initialization favors CP. Fig. 7(b) also shows the range of parameters over which we were able to find T solutions via matched asymptotics, showing that nonsymmetric initialization may not be sufficient to drive the system into T states, even if they exist. Together, these observations demonstrate multiple dynamical attractors, as may be expected for such complex PDE systems. Fig. S6 (in Appendix) shows an explicit example where both T and PS solutions exist for the same parameters. Understanding the basins of attraction of different states in more detail is an important challenge for future work.
Discussion
@@
We introduced the APLG as a microscopic model of interacting particles and characterized its hydrodynamic behavior. In addition to PS states familiar from pure active systems, we find rich behavior, including that the spinodal curve can protrude through the binodal. This signals the existence of dynamical steady states, which we classify according to their symmetries. Some of these results are similar to previous work on the NRCH equation [10, 9, 11, 13], but our hydrodynamic PDE includes the magnetization $m$ as a slow hydrodynamic variable, in addition to the two conserved densities $\rho_{a},\rho_{0}$; it also has a distinct set of nonlinearities. Hence, our approach of deriving hydrodynamic equations exactly from a microscopic model complements the generic description of nonreciprocal systems via the NRCH equation. As noted in Appendix .3, it is important that the APLG supports the complex behavior of other nonreciprocal models, despite the idealized modeling assumptions in its definition.
APLG phenomenology
@@
To tame the complexity of the APLG’s behavior, we focussed on large domains $L\gg 1$. This enables numerically exact computation of the binodal and spinodal curves and precise characterization of traveling solutions that involve sharp interfaces. Surprisingly, interfacial profiles in static and traveling states both obey (4), which also describes the tie-lines in the phase diagram. If such connections are generic in nonreciprocal systems, they would have broad consequences for understanding their phase diagrams, including possibilities for long-ranged order similar to equilibrium. We also demonstrated multistability: qualitatively different solutions can exist, for the same parameters.
@@
Our results also raise further questions for the APLG, including the existence of dynamical states in large systems with patterns on finite length scales, and associated questions of wavelength selection [13, 50, 11, 51]. While this work analyzed deterministic hydrodynamic equations, the theory of fluctuating hydrodynamics can also be developed for such models [18, 52, 53] by retaining corrections to the limit $h\to 0$. This enables studies of metastability and the role of noise in determining a system’s eventual steady state. There might also be interesting corrections to the large-$L$ behavior studied here. One may also expect new behavior on replacing the two-state active-particle orientation ($\sigma=\pm$) with a continuous degree of freedom: hydrodynamic limits can be derived in this case [19, 25] but the resulting equations are challenging to analyze. Future work should investigate these issues.
Implications for nonreciprocal systems
To provide a broader perspective, we note that the APLG model serves as an example of a general class of nonequilibrium mixtures. Its exact hydrodynamic limit relies on specific modeling assumptions, including a nontrivial $h$-dependence of the microscopic rates. However, the emergent large-scale picture is generic, including the appearance of complex stability eigenvalues, as observed in the NRCH and other nonreciprocal systems. As usual in mixtures of active and passive particles [22, 15, 9], this is driven by nonreciprocal effective interactions that appear in their hydrodynamic descriptions, notwithstanding that the microscopic interactions between particles are reciprocal.
For the APLG, we find that the protrusion of the spinodal through the binodal introduces a novel mechanism by which nonreciprocal behavior can arise, and it would be interesting to explore additional examples of this phenomenon. This could be achieved by considering hydrodynamic limits for other microscopic models, such as mixtures of different types of active particles [54]), or within the framework of field theories like the NRCH, independent of their underlying microscopic descriptions. For a recent example of an active system where the spinodal protrudes through the binodal, see [55], see also [56].
An important takeaway from this work is the tractability of the large-system limit. Large systems are essential for understanding equilibrium phase behavior, and it is valuable to identify nonequilibrium situations where such systems can be analyzed precisely. In the case of the APLG model, this is facilitated by the presence of domain walls, which appear both in static phase separation and in dynamical steady states. The scale separation between interfacial width and system size allows this behavior to be tackled using the method of matched asymptotics, providing a new route for characterizing dynamical pattern-forming states, such as those observed numerically in [22, 15, 54]. We look forward to future work in this direction, which holds promising opportunities to bridge theories of pattern formation with those of equilibrium phase behavior (see also Appendix B).
@@ Methods
For large systems, $L\gg 1$, stationary phase-separated (PS) states contain large domains of liquid and vapor phases. The phases’ (total) volume fractions are denoted $\phi_{l},\phi_{v}$, which we compute by generalising the method of [25, 57, 36]. Specifically, setting $\partial_{t}\rho=0=\partial_{t}\rho_{0}$ in (1) and integrating yields
@@
In the bulk of either phase, we have $\partial_{x}\rho_{\sigma}=0$, so from (11), $m=0$ there. Then evaluating Eqs. (9,10) in the bulk shows that $J=0=J_{0}$. Using this fact and eliminating $m$ between Eqs. (9,10), one obtains using (2) that $\partial_{x}\log[\rho_{0}/(1-\rho)]=0$ so $\rho_{0}=\nu(1-\rho)$ for some constant $\nu$. Integrating this equation over the whole domain, one obtains the value $\nu$ in (5). Then, combining Eqs. (4,9,11), we obtain a condition on $\rho(x)$ alone:
@@
with $g_{0},\Lambda,\kappa$ given in Eqs. (S32-S33).
The method of [25, 57, 36] can now be applied directly to (12). Within the bulk of the coexisting phases one has $\partial_{x}\rho=0=\partial_{x}^{2}\rho$, showing that $g_{0}(\phi_{l})=g_{0}(\phi_{v})$. In addition, (12) can be used to construct an effective free energy $\Phi$, from which $\phi_{l},\phi_{v}$ can be fully determined by a common tangent construction, see Appendix C for details. The compositions of the phases are then given by (4). A numerical implementation of this procedure yields the binodal curves in Fig. 1.
@@
To analyse the stability of homogeneous stationary solutions of (1), we introduce a perturbation to $\rho_{\sigma}$ constant of the form $\delta A^{\sigma}\exp(\lambda t+iqxt)$ for $\delta\ll 1$, and $\lambda$ can be obtained as the eigenvalue of a $3\times 3$ matrix. If $\text{Re}(\lambda)>0$, then the perturbation grows, signaling that the homogeneous state is unstable. The spinodal is the boundary between the regions of stable and unstable homogeneous solutions. Full details are given in Appendix D.
Traveling solutions via the Method of Matched Asymptotics.
To analyze T solutions in large domains, we define $\epsilon=1/L$ and seek a systematic approximation of (7) as $\epsilon\ll 1$ via the method of matched asymptotic expansions. Recall that $z\in[-L/2,L/2]$ and that, without loss of generality, the interface is centered at $z=0$ (see Fig. 6). We assume there is only one interface; the generalization to multiple interfaces is straightforward.
We define the outer region $|z|\gg 1$, where we set $z=\epsilon^{-1}\bar{z}$ and define $\varrho(z)=\bar{\varrho}(\bar{z})$ and $\mathrm{m}(z)=\bar{\mathrm{m}}(\bar{z})$ so (7) becomes
@@
with periodic boundary conditions at $\bar{z}=\pm 1/2$. Expanding $\bar{\varrho}_{\sigma}$ and $\bar{\mathrm{m}}$ in powers of $\epsilon$, $\bar{\varrho}_{\sigma}\sim\bar{\varrho}_{\sigma}^{(0)}(\bar{z})+\epsilon\bar{\varrho}_{\sigma}^{(1)}(\bar{z})+\cdots$ and $\bar{\mathrm{m}}\sim\bar{\mathrm{m}}^{(0)}+\epsilon\bar{\mathrm{m}}^{(1)}+\epsilon^{2}\bar{\mathrm{m}}^{(2)}+\cdots$, we find that the leading- and first-order of (14) lead to $\bar{\mathrm{m}}^{(0)}=0$ and $\bar{\mathrm{m}}^{(1)}=-(\text{Pe}/2)\partial_{\bar{z}}[d_{s}(\bar{\varrho}^{(0)})\bar{\varrho}_{a}^{(0)}]$, respectively. The $O(\epsilon^{2})$ of (14) is
$-c\bar{\varrho}_{\sigma}^{(0)\prime}=\partial_{\bar{z}}\left[d_{s}(\bar{\varrho}^{(0)})\bar{\varrho}_{\sigma}^{(0)\prime}+\bar{\varrho}_{\sigma}^{(0)}{\cal D}(\bar{\varrho}^{(0)})\bar{\varrho}^{(0)\prime}\right]-\sigma\bar{\mathrm{m}}^{(2)}\\ -\text{Pe}\partial_{\bar{z}}\!\left[\bar{\varrho}_{\sigma}^{(0)}s(\bar{\varrho}^{(0)})\bar{\mathrm{m}}^{(1)}\!+\!\sigma d_{s}(\bar{\varrho}^{(0)})\bar{\varrho}_{\sigma}^{(1)}\!+\!\sigma d_{s}^{\prime}(\bar{\varrho}^{(0)})\bar{\varrho}^{(1)}\bar{\varrho}_{\sigma}^{(0)}\right],$ (15)
using that $d_{s}(\bar{\varrho})\bar{\varrho}_{\sigma}\sim d_{s}(\bar{\varrho}^{(0)})\bar{\varrho}_{\sigma}^{(0)}+\epsilon d_{s}(\bar{\varrho}^{(0)})\bar{\varrho}_{\sigma}^{(1)}+\epsilon d_{s}^{\prime}(\bar{\varrho}^{(0)})\bar{\varrho}^{(1)}\bar{\varrho}_{\sigma}^{(0)}$ and $\bar{\mathrm{m}}^{(0)}=0$. Eliminating $\bar{\mathrm{m}}^{(2)}$ from (15) by adding them for $\sigma=\pm$ leads to
$-c\bar{\varrho}_{a}^{(0)\prime}=\partial_{\bar{z}}\left[d_{s}(\bar{\varrho}^{(0)})\bar{\varrho}_{a}^{(0)\prime}+\bar{\varrho}_{a}^{(0)}{\cal D}(\bar{\varrho}^{(0)})\bar{\varrho}^{(0)\prime}\right]\\ -\text{Pe}\partial_{\bar{z}}\left[\left(\bar{\varrho}_{a}^{(0)}s(\bar{\varrho}^{(0)})+d_{s}(\bar{\varrho}^{(0)})\right)\bar{\mathrm{m}}^{(1)}\right],$ (16)
using again that $\bar{\mathrm{m}}^{(0)}=0$. Inserting the expression for $\bar{\mathrm{m}}^{(1)}$ into Eqs. (15, 16) and switching back to $z$ leads to (8) at leading order in $\epsilon$.
To solve (8), it remains to determine the boundary conditions as we approach the interface, or $|\bar{z}|\to 0$. In the inner region $z=O(1)$, define densities $\hat{\varrho}_{\sigma}(z)=\varrho_{\sigma}(z)$. They solve (7), together with the matching condition to the outer region, $\lim_{z\to\pm\infty}\hat{\varrho}(z)=\lim_{\bar{z}\to 0^{\pm}}\bar{\varrho}(\bar{z})$; these limits exist under the assumptions of a localized interface. At leading order in $\epsilon$, the LHS of (7) vanishes: this ensures consistency with the argument above that traveling phase-separated states do not exist, and justifies the scaling of the speed as $c/L$ in (7). Hence, the leading-order inner solution solves the same equations as the PS state [Eqs. (9,10,11) with $J=0=J_{0}$], and interfaces in T states connect points on the binodal curve.
@@
$\bullet$ Particle model: The APLG is a continuous time Markov chain on a finite state space. We simulate it exactly with the Gillespie algorithm [40], initially placing a $\sigma$-particle on a lattice site with probability $\phi_{\sigma}$, where $\phi_{\pm}=\phi_{a}/2,\phi_{0}=\phi_{p}$. In Figs. 2,3, the simulated domain has $\ell_{y}=\ell_{x}/4$; we plot the $y-$averaged values of the mesoscopic densities (see (S2)) with radius $r=0.1$.
$\bullet$ Time-stepping of Eq. (1): We use a first-order finite-volume scheme in space and forward Euler with adaptive time-stepping in time to obtain time-dependent numerical solutions $\rho_{\sigma}(x,t)$ to the hydrodynamic PDE (1), building on the numerical scheme of [25, 43] (see Appendix 1).
$\bullet$ Classification of the steady-states: We classify solutions of (1) into Homogeneous (H), Phase-separated (PS), Traveling (T), or Counter-propagating (CP) using the following two metrics: the approximate speed $\tilde{c}(t):=\|\partial_{t}\rho_{\sigma}\|_{2}/\|\partial_{x}\rho_{\sigma}\|_{2}$ and the distance from uniform
@@
We solve (1) until a final time $t^{*}\geq 700$ when one of the following conditions is satisfied (where $\mathcal{T}^{*}=[t^{*}-500,t^{*}]$):
@@
    $d_{\text{H}}(t^{*})\geq 0.05$ and $\sup_{t\in\mathcal{T}^{*}}\tilde{c}(t)<0.01$to PS solution.
@@
    $d_{\text{H}}(t^{*})\geq 0.05$, $\tilde{c}(t^{*})\geq 0.01$ and $\sup_{t\in\mathcal{T}^{*}}|\tilde{c}^{\prime}(t)|<10^{-5}$to T solution.